OpenStudy (anonymous):

The length of each side of a square is extended 4 cm. The area of the resulting square is 100 cm2. Find the length of a side of the original square.

OpenStudy (anonymous):

Let the original length be x (x+4)^2 = x^2+8x+16 = 100 x^2 +8x - 84 (x-6)(x+14) x = 6, -14